How to plot graphically the savings function derived algebraically, indicating the positions Y>C, Y<C, Y=C


1
Expert's answer
2021-05-18T12:14:32-0400

From the keynes consumption and income function it is known that ,

"Y=C+S \\

C=a+MPC(Y)\\

S=Y-C\\

S=-a+Y(1-MPC)"

Where a is the a= autonomous consumption function.

And savings is a positive function of Income therefore positive and upward sloping.


Now, considering three situations, when Y=C, there will be no savings as at point Y0, when Y>C at Y2 it means that there is some savings which is denoted by a. When Y<C at point Y1 then it means the consumer borrows and the exogenous consumption becomes -a.